An economy is described by the following equation
Cd=14400+0.5(Y-T)-40000r, Ip=8000-20000r,
G=7800, NX=1800, T=8000
a) Find the numerical equation relating planned aggregate expenditure (PAE) to output (Y) and to real interest rate (r).
b) The real interest rate is 0.133, find short-run equilibrium output.
c) Potential output, y*, equals 40,000. What real interest rate should be Reserve Bank set to bring the economy to full employment?
Answer #1

Planned Aggregate Expenditure (PAE) is given as: PAE = Cd +Ip+G+NX =14.400 +0.5(Y -8.000)-40.000r+17.600 - 20.000r = 32.000 +